40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 18 18:33:57 2014, ericq, Update, Green Locking, Yarm Green PDH yLoop.pdfcrap.pdf
    Reply  Tue Aug 19 01:01:36 2014, Jenne, Rana, Update, Green Locking, Yarm Green PDH YPDH_OLG.pdfYPDH_noise.pdf
       Reply  Tue Aug 19 18:32:40 2014, ericq, Update, Green Locking, Yarm Green PDH 
       Reply  Tue Aug 19 23:11:15 2014, Jenne, Update, Green Locking, Yarm Green PDH ErrAndCtrlSpectra_VoltsPerRtHz.png
       Reply  Wed Aug 20 02:38:41 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement NoiseConsideration.pdf
          Reply  Wed Aug 20 18:05:18 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement YPDH_noise.pdf
       Reply  Wed Aug 20 04:09:21 2014, ericq, Update, Green Locking, Xarm Green PDH Xbodes.pdfXspectra.pdfscopeSweep.jpg
          Reply  Wed Aug 20 16:10:43 2014, ericq, Update, Green Locking, Xarm Green PDH XspectraCombined.pdf
             Reply  Wed Aug 20 21:09:16 2014, ericq, Update, Green Locking, Xarm Green PDH scopeSweep.jpgXspectra.pdfYspectra.pdfXYcomp.pdf
Message ID: 10407     Entry time: Mon Aug 18 18:33:57 2014     Reply to this: 10408
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: Yarm Green PDH 

So far today, I've been working with the Y-end green PDH locking. Using a SR560 to roll off the AG4395A output to take a loop measurement at the servo output, I measured the following OLG, and inferred the CLG from it. The SR560 really helped it getting good coherence without introducing a big offset that changes the optical gain, thus distorting the loop shape, etc. etc. 

yLoop.pdf

You would think this loop looks pretty good, 10k UGF, and 45 degrees of phase margin, gain peaking is sane, and pretty smooth slope. But, the thing still was flipping out of lock while I measured this. 

I suspect shenanigans at >100k. This is motivated by the fact that I've seen some big noise in the error signal around 150k. I don't have a good noise plot right now, because I'm trying to get a scheme going where I stitch together a bunch of 1 decade spectra from the 4395, but the noise floor isn't consistent across each patch (even though the attenuation stays the same, and I confirmed I'm in "noise" mode). I'm working on a loop measurement up there, too, but I haven't been able to get the right filter/amplitude settings yet. 

So, even though this plot is not totally correct (read: wrong and bad), I include it just for the sake of showing the big honking spike of noise at ~150K.  

crap.pdf

 

ELOG V3.1.3-